# Explainable Visual Question Answering with Multimodal Rationales

### v2

@@ III-C3 Projection Module
where $C$ represents the learnable constants, q, K, V and d denote query, Key, Value, and dimensionality of values, respectively. We fuse the projected features $I_{proj}$ and $Q_{proj}$ by concatenating before feeding the results $f_{proj}$ into the LLMs.
III-C4 Predictors
Since the multimodal EVQA task encompasses generating answers, textual rationales, and visual rationales, we utilize different predictors in our model. For the answer prediction, we approach it as a classification problem, similar to previous methods [2, 19]. Specifically, we utilize a linear fusion module, followed by a sigmoid function, to process the cross-attended features from the decoder. For textual rationales, in contrast to earlier methods that rely on simple LSTMs [29], we utilize advanced LLMs to generate more robust and relevant textual rationales. However, given the complexity of handling multiple tasks in EVQA, the LLMs integrated into the model should facilitate computationally efficient fine-tuning for multimodal learning. To address this, we leverage the auto-regressive large language model GPT-2 [41]. For each token in the sequence, GPT-2, conditioned on the projected features $f_{proj}$, generates probability distributions over the entire vocabulary. These distributions are then used to predict the next token in the sequence, enabling the generation of complete textual rationales. For visual rationales, we generate precise object-level bounding boxes using Grounding-DINO [1], an open-vocabulary object detector capable of identifying arbitrary objects based on natural language inputs. In contrast to the original work, our approach inputs the detector with the image $I$, cross-attended features $f_{cross}$, and LLM-processed features $f_{llm}$. In addition, we set all detected bounding boxes as a single category, diverging from standard visual grounding practices [32]. This setting addresses the challenge of aligning objects with the input question when specific descriptions are absent. For example, when given the phrase “a red baseball bat”, visual grounding models can effectively match the object “bat” to the relevant image regions. However, for a question like “What color is the object?”, the models might struggle to discern the category information from the image-question pair due to the lack of explicit context. With the incorporation of various predictors, our model provides robust results for multimodal EVQA through an end-to-end supervised mode.
III-C5 Loss Functions
During model training, we use full supervision with annotated examples to generate answers, textual rationales, and bounding boxes. A hybrid loss function is employed to refine the model, with different loss components tailored for each type of prediction in our EVQA task. For the answer prediction, we use a sigmoid function to output a score $\hat{s}$ ranging from 0 to 1, representing the probability of the answer. Consistent with previous methods [9], we employ the widely-used binary cross-entropy loss [42] as $L_{ans}$ to supervise these predictions. To train the model for generating textual rationales from the LLM-processed features $f_{llm}$, we utilize the simple yet effective cross-entropy loss as $L_{tr}$ in our case. Given $M$ questions, $N$ candidate answers, $M$ textual rationales, $L_{ans}$ and $L_{tr}$ are calculated as follows,
@@ III-C5 Loss Functions
where $l$ and $\theta$ represent the maximum length of tokens and the model’s trainable parameters, respectively. The sequenced tokens are denoted as $w^{i}=w_{1}^{i},...,w_{l}^{i}$. For visual rationale generation, we directly adopt the existing loss function from Grounding-DINO [1], including L1 loss, GIoU loss, and contrastive loss. We combine them together as our vision objective $L_{vr}$ for brevity. To train the entire model using multi-task learning, we take a linear combination of the different loss components as follows,
$L_{base}=L_{ans}+L_{tr}+L_{vr},$ (6)
where $L_{base}$ denotes the base objective of our multimodal EVQA task. Note that we remove the final component of $L_{base}$ for our MRVQA-C model due to the lack of visual supervision from the annotations.
@@
$L=L_{base}+\lambda L_{ta},$ (8)
Compared to previous single-track and compositional EVQA methods [9, 13], our proposed loss function ensures greater consistency between answer prediction and textual rationale generation, avoiding the disconnection often seen with multiple predictors. In addition, the visual rationale generation leverages features from both the answer and text spaces, contributing to a unified model.
III-D Visual-Textual Similarity Score
Previous methods for text generation tasks (e.g., translation and retrieval) commonly utilize word-level or phrase-level similarity evaluation metrics, such as BLEU [14], ROUGE [15] and SPICE [43], to evaluate model performance. However, these metrics primarily capture low-level aspects (e.g., fluency and coherence), while overlooking the high-level interactions between sentences that do not strictly match word-for-word. Fortunately, the similarity score between sentence embeddings is a good option to measure the quality of generated texts. We utilize the powerful pre-trained text embedding model, GTE [44], to map the generated textual rationales for similarity comparison. There are several reasons for selecting this model: (1) The GTE model employs contrastive learning across a diverse range of datasets, which allows it to generalize effectively to various data; (2) It avoids the reliance on prompt formulation seen in LLMs; (3) It handles pairs of short and long texts more effectively without requiring word-level alignments, which is ideal for comparing our generated free-form texts; (4) It outperforms previous embedding methods in comprehensive benchmarks [44]. We believe the GTE model is well-suited to measure the similarity between the generated textual rationales and the ground truth. Specifically, we use the cosine function to assess textual similarity. To handle the negative values produced by the cosine function, we map the cosine similarity scores to a range from 0 to 1. The textual similarity ($TS$) is calculated as follows,
@@ III-D Visual-Textual Similarity Score
where $T_{pred}$ and $GT$ represent the generated textual rationale and the ground truth, respectively. For visual evaluation, we assess the predicted bounding boxes by calculating the Average Precision [45] (AP) in comparison to ground truths.
For a unified multimodal model, it is essential to evaluate its performance from both visual and textual perspectives, ensuring coherence across different predictions. However, simply averaging TS and AP scores fails to fully capture the model’s performance, particularly when one modality significantly lags behind the other in certain scenarios (e.g., occlusion or ambiguity). To address this, we introduce a new metric vtS (visual-textual Similarity score), which combines visual evaluation and textual similarity to provide a balanced assessment of the quality of the generated multimodal rationales. The metric vtS is defined as follows,
@@
Given the interdependence between visual and textual rationales, more detailed and relevant texts can improve the extraction of bounding boxes. Likewise, more accurate visual rationales help guide the generation of more focused and precise textual descriptions. The combined evaluation metric provides a unified way to assess overall model performance.
IV Experiments
IV-A Experiment Setup
IV-A1 Datasets
We conduct the primary experiments on the synthesized multimodal EVQA dataset (VQA-E-Syn) introduced in Section III-B. In addition, we evaluate models on three existing VQA datasets, including VQA-X [8], GQA-REX [13], and VQAv2 [46], to access the its adaptability and generalization capability under different settings.
IV-A2 Training and Inference Settings
Under full supervision using the synthesized dataset, we refine the MRVQA model in an end-to-end manner, referred to as MRVQA-E. The proposed loss function is designed to help optimize the training process, enhancing the model’s ability to generate both answers and multimodal rationales. To adapt to various practical scenarios, we simplify the MRVQA model to predict the required results under different conditions. For example, we develop a variant, MRVQA-C (combined), by removing the visual generator component, allowing us to test the model’s ability to generalize in the absence of visual rationale supervision. This variant utilizes the generated textual rationales and the input images to guide the Grounding-DINO [1] model in predicting bounding boxes during inference stage. In addition, we further simplify the MRVQA model to focus exclusively on answer prediction or textual rationale prediction, developing variants called MRVQA-A and MRVQA-TR for direct comparison with previous standard VQA methods and single-track EVQA methods, respectively. Note that only the relevant loss functions are utilized for these simplified models.
TABLE II: Comparison results on VQA-E-Syn regarding textual rationale generation. All numbers are reported in percentage (%). The best results are bold while the second best are underlined.
@@ IV-A2 Training and Inference Settings
MRVQA-TR CLIP CLIP GPT2 15.08 21.77 42.04 102.86 21.22 57.30
MRVQA-C CLIP CLIP GPT2 15.84 22.41 43.57 105.31 22.19 58.19
MRVQA-E CLIP CLIP GPT2 16.97 23.02 44.28 107.04 23.68 59.16
IV-A3 Implementation Details
We utilize the pre-trained CLIP model [38] with ViT-base [39] and Transformer [17] architectures for visual and linguistic feature representation. The cross-attended and LLM-processed features are provided as additional inputs to the feature enhancer of the Grounding-DINO-T model. We use the AdamW [49] optimizer with the learning rate of $2\times 10^{-5}$. All the models are trained for 50 epochs with a batch size of 64 using two A40 GPUs, within a consistent PyTorch environment. To ensure a fair comparison, we follow the official repositories of the relevant methods and reimplemented all the models on the synthesized dataset.
IV-A4 Evaluation Metrics
For VQA evaluation, we maintain consistency with prior works by using answer accuracy as the metric. To evaluate generated textual rationales, we employ five mainstream metrics, namely BLEU-4 [14], METEOR [50], ROUGE [15], CIDEr [16] and SPICE [43] (B, M, R, C, S for their abbreviations, respectively), which measures the linguistic similarity between the generated text and the ground truth. Additionally, we use the newly introduced metric vtS to better capture the cross-modality similarity by leveraging the visual response. The widely-used Average Precision (AP) metric [45] is employed to assess the quality of the generated bounding boxes by comparing them to ground truths.
IV-B Comparisons with the State-of-the-Art Methods
IV-B1 Results on the synthesized VQA-E-Syn
Table II reports the comparison with the state-of-the-art EVQA methods on the synthesized VQA-E-Syn dataset. Our proposed MRVQA-E model significantly enhances the generation of textual rationales for multimodal EVQA, outperforming the previous methods. The achieved higher scores on the evaluation metrics demonstrate that our generated text excels in fluency, relevance, and accuracy. For example, our approach achieves a 2.41% improvement in CIDEr compared to VCIN [31], highlighting its superior fluency and coherence. For evaluating multimodal rationales, we report results using the introduced vtS. It is worth mentioning that we employ a combined approach similar to our MRVQA-C variant to gather visual predictions for comparison with the previous methods that do not support that. The results demonstrate that our approach achieves a 1.27% improvement over VCIN in vtS. To ensure a fair comparison, we also evaluate our MRVQA-TR and MRVQA-C models. Both variants demonstrate slightly superior performance compared to the previous methods. More importantly, the comparison between the end-to-end model and its variants highlights that incorporating visual rationale generation improves linguistic prediction performance.
TABLE III: Comparison results on VQA-E-Syn regarding answer accuracy. OA and AA represent overall and average accuracy, respectively.
@@ IV-B1 Results on the synthesized VQA-E-Syn
MRVQA-A 57.02 72.80 73.41 71.89 67.75
MRVQA-C 63.20 76.03 75.72 74.81 71.65
MRVQA-E 65.11 76.40 75.87 75.01 72.46
Table III shows the results compared to the state-of-the-art VQA methods on VQA-E-Syn in terms of answer accuracy. The methods listed in the top section are trained solely for answer prediction, whereas those in the middle section have the capability to generate both answers and textual explanations. For example, MCAN [23], LXMERT [51], and UNITER [52] are popular models for standard VQA. OFA [48], as a powerful unified sequence-to-sequence framework, facilitates generating multiple predictions through separate fine-tuning processes on task-specific datasets (e.g., VQA and image captioning). To adapt OFA for our EVQA task, generating texts within a unified inference pass, we perform a single fine-tuning of the OFA-base model on the combined samples from VQA-E-Syn. Our MRVQA-E model achieves an overall accuracy of 75.01% and an average accuracy of 72.46% on VQA-E-Syn, surpassing the previous methods. Notably, the model shows significant improvements in the Number and Yes/No question categories, with increases of 1.43% and 2.59%, respectively. Similarly, comparing MRVQA-E with its variants (last three rows) reveals that incorporating additional rationale supervision enhances the performance of the answering process. This approach provides more detailed information, significantly benefiting the VQA model.
IV-B2 Results on VQA-X
Table IV reports the results compared to previous EVQA methods on VQA-X [8]. As VQA-X only supports linguistic examples (shown in Table I), to ensure a fair comparison, we conduct experiments with MRVQA-C for only answer and textual rationale generation. Our approach achieves a 78.8% overall accuracy, which is higher than the previous state-of-the-art VCIN by 1.1%. The comparison of the generated textual explanations across various metrics also demonstrates that our approach surpasses previous methods. For instance, the model shows significant improvements in BLEU-4 and CIDEr metrics, with increases of 0.7% and 4.7%, respectively, compared to VCIN.
TABLE IV: Comparison results on VQA-X [8] and GQA-REX [13] for answer and textual rationale evaluation.
@@ IV-B2 Results on VQA-X
IV-B3 Results on GQA-REX
Table IV also shows the results on GQA-REX [13]. Note that the previous VCIN model predicted region-level patches and compared results with other adapted methods using the Grounding metric [13], which relied on priors in visual encodings (e.g., patch location). Due to the single-category object detection used in our settings, we exclude this unfair comparison and only report the textual results. Our approach achieves an overall answer accuracy of 61.28%, representing a 0.67% improvement over the previous state-of-the-art VCIN. Meanwhile, our model maintains competitive performance even without external information (e.g., functional program and scene graph), achieving 58.83% and 83.07% for BLEU-4 and ROUGE scores, respectively. These results highlight the model’s strong generalization capability across different datasets.
IV-B4 Results on VQAv2
We further assess our model’s transfer learning capabilities on VQAv2 [46] and compare the results with previous methods in Table V. The top section of the table displays methods that were either trained from scratch or pre-trained on large-scale datasets before being fine-tuned on VQAv2. As expected, results improve with the use of more powerful models and larger datasets. For example, BLIP [4], which is pre-trained on over 129 million image-text pairs, achieves 78.3% accuracy after thorough fine-tuning on VQAv2. The methods listed in the lower section of the table are trained on alternative datasets without subsequent fine-tuning, allowing us to assess the robustness and generalization capability of these models. For example, we directly evaluate our pre-trained MRVQA-E model on VQAv2. Despite being trained on relatively limited data (23K) compared to other methods, our approach achieves competitive performance, reaching 66.8% accuracy on a standard VQA task. The comparison with BLIP2 [54] highlights the effectiveness of integrating rationales into the answering system.
[…]
Method Training Dataset Data Scale Fine-tuned? OA
[…]
IV-C Ablation Studies
IV-C1 Ablation of the Projection Module
In the proposed MRVQA model, we introduce a Transformer-based projection module to map the pre-trained CLIP features to the latent space of LLMs. For comparison, we conduct an ablation study by replacing the projection module in MRVQA-E with a linear layer, as used in prior vision-language methods. The results in Table VI show that our Transformer-based projection module significantly enhances the performance of textual rationale generation compared to the linear layer across various metrics. For example, the incorporation of our module results in a 5.46% improvement in CIDEr score. Additionally, the enhancement in linguistic predictions contributes to a 2.02% increase in the multimodal rationale evaluation metric vtS, highlighting the efficacy of our proposed module.
TABLE VI: Ablation results of using different projection modules for multimodal rationale generation on VQA-E-Syn.
@@ IV-C1 Ablation of the Projection Module
IV-C2 Ablation of Unified or Separate Text Predictor
For text predictions, we utilize two separate predictors: a classifier for answer prediction and a generator for textual rationale prediction. This approach is designed to ensure high accuracy in answer predictions with a concise format (i.e., short and straightforward), making the results more accessible to non-expert users. Additionally, this separation avoids the need for redundant task-specific fine-tuning, as seen in previous pre-training methods [52, 48]. Some approaches [54] use a unified strategy for all text-related downstream tasks, which often results in longer and more complex expressions. To compare it with our approach, we adapt the unified strategy by integrating answer prediction into the text generation process, using samples in {answer, textual rationale} format (e.g., “the boy because only the boy can catch the frisbee”). Since the primary focus of this ablation is on the text prediction, we use the MRVQA-C model to ensure a fair comparison. Table VII reports the performance of different text prediction strategies. As expected, while the performance of textual rationale generation shows only a slight change, the overall accuracy significantly drops by 4.39%, indicating a decrease in the robustness of an answering model. Consequently, we employ dual-head predictors for linguistic output to address this issue in our EVQA task.
TABLE VII: Ablation results of different strategies for text predictions on VQA-E-Syn.

### v3

@@ 3.3 Multimodal EVQA Network
where $C$ represents the learnable constants, q, K, V and d denote query, Key, Value, and dimensionality of values, respectively. The projected features $I_{proj}$ and $Q_{proj}$ are fused by concatenation as $f_{proj}$ before fed into the LLMs.
Predictors. For the answer prediction, we approach it as a classification problem and utilize a linear fusion module, followed by a sigmoid function, to process the cross-attended features from the decoder. For textual rationales, in contrast to earlier methods that rely on simple LSTMs [15], we utilize advanced LLMs to generate more robust and relevant textual rationales. However, given the complexity of handling multiple tasks in EVQA, the LLMs integrated into the model should facilitate computationally efficient fine-tuning for multimodal learning. To address this, we leverage the auto-regressive large language model GPT-2 [38]. For each token in the sequence, GPT-2, conditioned on the projected features $f_{proj}$, generates probability distributions over the entire vocabulary and predicts the next token in the sequence. For visual rationales, we generate precise object-level bounding boxes using Grounding-DINO [29], an open-vocabulary object detector based on text. In contrast to the original work, our approach inputs the detector with the image $I$, cross-attended features $f_{cross}$, and LLM-processed features $f_{llm}$. In addition, we set all detected bounding boxes as a single category. This setting addresses the challenge of aligning objects with the input question for discerning the category information from the image-question pair due to the lack of explicit context.
Loss Functions. A hybrid loss function is employed to refine the proposed model, with different loss components tailored for each type of prediction in our EVQA task. Consistent with previous methods [23], we employ the widely-used binary cross-entropy loss [42] as $L_{ans}$ to supervise answer predictions. To train the model for generating textual rationales from the LLM-processed features $f_{llm}$, we utilize the simple yet effective cross-entropy loss as $L_{tr}$ in our case. Given $M$ questions, $N$ candidate answers, $M$ textual rationales, $L_{ans}$ and $L_{tr}$ are calculated as follows,
@@
where $l$ and $\theta$ represent the maximum length of tokens and the model’s trainable parameters, respectively. The sequenced tokens are denoted as $w^{i}=w_{1}^{i},...,w_{l}^{i}$. For visual rationale generation, we directly adopt the existing loss function from Grounding-DINO [29] as $L_{vr}$, including L1 loss, GIoU loss, and contrastive loss. To train the entire model using multi-task learning, we take a linear combination of the different loss components.
@@
$L=L_{ans}+L_{tr}+L_{vr}+\lambda L_{ta},$ (7)
3.4 Visual-Textual Similarity Score
We utilize the powerful pre-trained text embedding model, GTE [25], to map the generated textual rationales for similarity comparison to the ground truths. There are several reasons for selecting this model: (1) The GTE model employs contrastive learning across a diverse range of datasets, which allows it to generalize effectively to various data; (2) It avoids the reliance on prompt formulation seen in LLMs; (3) It handles pairs of short and long texts more effectively without requiring word-level alignments, which is ideal for comparing our generated free-form texts; (4) It outperforms previous embedding methods in comprehensive benchmarks [25]. Specifically, we use the cosine function to assess textual similarity. To handle the negative values produced by the cosine function, we map the cosine similarity scores to a range from 0 to 1. The textual similarity ($TS$) is calculated as follows,
@@ 3.4 Visual-Textual Similarity Score
where $T_{pred}$ and $GT$ represent the generated textual rationale and the ground truth, respectively. We assess the predicted bounding boxes by calculating the Average Precision [14] (AP) in comparison to ground truths.
To this end, we introduce a new metric vtS (visual-textual Similarity score), which combines visual evaluation and textual similarity to provide a balanced assessment of the generated rationale quality. The metric vtS is defined as follows,
@@
4 Experiments
4.1 Experiment Setup
Datasets. We conducted primary experiments on the synthesized multimodal EVQA dataset (VQA-E-Syn) introduced in Section 3.2. In addition, we evaluated the model on two existing EVQA datasets: VQA-X [34] and GQA-REX [8], to access the its adaptability and generalization capability under different settings.
Training and Inference Settings. Under full supervision, we refined the model in an end-to-end manner, referred to as MRVQA-E. To adapt to various practical scenarios, we simplified the MRVQA model to predict the required results under different conditions. For example, we developed a variant, MRVQA-C (combined), by removing the visual generator component, allowing us to test the model’s ability to generalize in the absence of visual rationale supervision. This variant utilized the generated textual rationales and input images to guide the detector in predicting bounding boxes during inference stage. In addition, we further simplified the model to focus exclusively on answer prediction or textual rationale prediction, developing variants called MRVQA-A and MRVQA-TR, respectively.
Table 1: Comparison results on VQA-E-Syn regarding textual rationale generation. All numbers are reported in percentage (%). The best results are bold while the second best are underlined.
@@ 4.1 Experiment Setup
MRVQA-TR (ours) CLIP CLIP GPT2 15.08 21.77 42.04 102.86 21.22 57.30
MRVQA-C (ours) CLIP CLIP GPT2 15.84 22.41 43.57 105.31 22.19 58.19
MRVQA-E (ours) CLIP CLIP GPT2 16.97 23.02 44.28 107.04 23.68 59.16
Implementation Details. We utilized the pre-trained CLIP model [39] with ViT-base [12] and Transformer [43] for input representation. The AdamW [18] optimizer was used with a learning rate of $2\times 10^{-5}$. All the models were trained for 50 epochs with a batch size of 64 using two A40 GPUs, within a consistent PyTorch environment.
Evaluation Metrics. For answer evaluation, we maintained consistency with prior works by using accuracy as the metric. To evaluate generated textual rationales, we employed five mainstream metrics, namely BLEU-4 [33], METEOR [4], ROUGE [26], CIDEr [44] and SPICE [1] (B, M, R, C, S for their abbreviations, respectively). Additionally, the newly introduced metric vtS was used to capture the cross-modality similarity by leveraging the visual response.
4.2 Quantitative Results and Comparison
4.2.1 Results on the synthesized VQA-E-Syn
Table 1 presents a comparison with state-of-the-art EVQA methods trained and evaluated on the synthesized VQA-E-Syn dataset. Our proposed MRVQA-E model significantly enhanced the generation of textual rationales for multimodal EVQA, outperforming the previous methods. For example, our approach achieved a 2.41% improvement in CIDEr compared to VCIN [47], highlighting its superior fluency and coherence. To ensure a fair comparison with previous methods lacking visual prediction capabilities, we adopted a combined approach similar to our MRVQA-C variant. The results demonstrate that our approach achieved a 1.27% improvement over VCIN in vtS. Additionally, we also evaluated our MRVQA-TR and MRVQA-C models. Both variants demonstrated slightly superior performance compared to the previous methods. More importantly, the comparison between the end-to-end model and its variants highlights that incorporating visual rationale generation improves linguistic prediction performance.
Table 2: Comparison results on VQA-E-Syn regarding answer accuracy. OA and AA represent overall and average accuracy, respectively.
@@ 4.2.1 Results on the synthesized VQA-E-Syn
MRVQA-A (ours) 57.02 72.80 73.41 71.89 67.75
MRVQA-C (ours) 63.20 76.03 75.72 74.81 71.65
MRVQA-E (ours) 65.11 76.40 75.87 75.01 72.46
Table 2 shows the results in terms of answer accuracy. The methods listed in the top section are trained solely for answer prediction, whereas those in the middle section have the capability to generate both answers and textual explanations. OFA [45], as a powerful unified sequence-to-sequence framework, facilitates generating multiple predictions through separate fine-tuning processes on task-specific datasets (e.g., VQA and image captioning). To adapt OFA for our task, generating texts within a unified inference pass, we performed a single fine-tuning of the OFA-base model on the combined samples from VQA-E-Syn. Methods lacking answer prediction details or accuracy results [34, 46, 35] are excluded from this table. Our MRVQA-E model achieved an overall accuracy of 75.01% and an average accuracy of 72.46% on VQA-E-Syn, surpassing the previous methods. Notably, the model showed significant improvements in the Number and Yes/No question categories, with increases of 1.43% and 2.59%, respectively. Similarly, comparing MRVQA-E with its variants (last three rows) reveals that incorporating additional rationale supervision enhanced the performance of the answering process.
4.2.2 Results on VQA-X
Table 3 reports the results compared to previous EVQA methods on VQA-X [34]. As VQA-X only supports linguistic examples, to ensure a fair comparison, we conducted experiments with MRVQA-C for only answer and textual rationale generation. Our approach achieved a 78.8% overall accuracy, which is higher than the previous state-of-the-art VCIN by 1.1%. Also, the model showed significant improvements in BLEU-4 and CIDEr metrics, with increases of 0.7% and 4.7%, respectively, compared to VCIN.
Table 3: Comparison results on VQA-X [34] and GQA-REX [8] for answer and textual rationale evaluation.
@@ 4.2.2 Results on VQA-X
4.2.3 Results on GQA-REX
Table 3 also shows the results on GQA-REX [8]. Note that the previous VCIN model predicts region-level patches, which relies on priors in visual encodings (e.g., patch location). Due to the single-category object detection used in our settings, we exclude this unfair comparison and only report the textual results. Our approach achieved an overall answer accuracy of 61.28%, representing a 0.67% improvement over the previous state-of-the-art VCIN. Meanwhile, our model maintained competitive performance even without external information, achieving 58.83% and 83.07% for BLEU-4 and ROUGE scores, respectively.
4.3 Ablation Studies
4.3.1 Ablation of the Projection Module
For comparison, we conducted an ablation study by replacing the projection module in MRVQA-E with a linear layer, as used in prior vision-language methods. The results in Table 4 show that our Transformer-based projection module significantly enhanced the performance of textual rationale generation compared to the linear layer across various metrics. For example, the incorporation of our module resulted in a 5.46% improvement in CIDEr score. Additionally, the enhancement in linguistic predictions contributed to a 2.02% increase in the multimodal rationale evaluation metric vtS, highlighting the efficacy of the module.
Table 4: Ablation results of using different projection modules for multimodal rationale generation on VQA-E-Syn.
@@ 4.3.1 Ablation of the Projection Module
4.3.2 Ablation of Unified or Separate Text Predictor
Some approaches [21] use a unified strategy for all text-related downstream tasks, which often results in longer and more complex expressions. To compare it with our approach, we adapt the unified strategy by integrating answer prediction into the text generation, using samples in {answer, textual rationale} format (e.g., “the boy because only the boy can catch the frisbee”). Table 5 reports the performance of different text prediction strategies. While the performance of textual rationale generation showed a slight change, the overall accuracy significantly dropped by 4.39%, indicating a decrease in the robustness of an answering model. Thus, we employed dual-head predictors for linguistic outputs to address this issue in the EVQA task.
Table 5: Ablation results of different strategies for text predictions on VQA-E-Syn.